# A novel two-field pure K-essence for inflation, dark matter, dark energy and black holes

### 06 Feb 2024 v1

@@ I Introduction
K-essence is a certain class of scalar field models with nonstandard kinetic terms. The name was first coined in a paradigm for inflation by Armendariz-Picon et al. picon:1999 ; garriga:1999 . Subsequently, it was found that k-essence can present interesting models for dark energy chiba:2000 ; picon:2000 ; picon:2001 ; chiba:2002 ; chimento:2004 ; chimento:2003 ; malquarti:2003 . Right after these studies, Scherrer shows that K-essence can also serve as the unified model for dark matter and dark energy robert:2004 . Afterwards, researches on the K-essence theories spring up.
@@
Specifically, Chimonto and Lazkoz make an analysis on the atypical K-essence cosmologies chimento:2005 . Abramo et al. conduct the research on the stability problem of phantom k-essence theories abramo:2006 . Bonvin et al. bonvin:2006 demonstrate that if one use K-essence to solve the coincidence problem and play the role of dark energy in the Universe, the fluctuations of the field have to propagate superluminally at some stage. This implies that successful K-essence models violate causality. Therefore, k-essence cannot arise as a low energy effective field theory of a causal, consistent high energy theory. Bamba bamba:2012 et al carry out a thorough investigation on the cosmological perturbations of K-essence. The causality and superluminal behavior K-essence theories are investigated by Bruneton jean:2007 . The slow-roll conditions for thawing K-essence are derived by Chiba et al. with a separable Lagrangian K-essence chiba:2009 . Considering the unification between the swampland criteria and the attractor, Herrera studies the reconstruction of the K-essence inflationary universe herrera:2020 . Furthermore, K-essence can yield a complete picture of the evolution of the Universe, starting from the early inflation, the subsequent dark matter domination, and to the late time cosmic acceleration bose:2009a ; bose:2009b . Interestingly, Tian and Zhu’s research tian:2021 reveals that k-essence framework is also conducive to the Hubble tension problem riess:2021 ; freedman:2021 .
@@
Very recently, Hung and Miao hung:2023 investigate whether a quantum-induced k-essence can provide efficient reheating without affecting the observational constraints on primordial inflation. Kehayias and Scherrer find a new generic evolution for K-essence dark energy with the equation of state (EOS) of roughly cosmological constant kehayias:2019 . Sarkar et al sarkar:2023 show that the accretion of k-essence dark energy can lead to a faster circularising of orbits for binary black holes. Lara et al lara:2022 investigate in depth the relation between K-essence, well-posedness of the Cauchy problem and the UV completions. Huang huang:2021 constructs the thawing K-essence models by generating Taylor expansion coefficients of Lagrangian functions from random matrices. Chatterjee et al. make a study on the dynamical stability problem for k-essence interacting nonminimally with a perfect fluid chatt:2021 . Applying the “$2+1+1$” decomposition of spacetime, Nagy et al conduct the research of spherically symmetric, static black holes with scalar hair, and naked singularities in nonminimally coupled K-essence nagy:2021 . Barvinsky et al. find that the generalized unimodular gravity can be transformed into a new form of K-essence barvinsky:2021 . Finally, the K-essence model in which K-essence can be responsible for both primordial inflation and the present observed acceleration, while also admitting a nonsingular de Sitter beginning of the Universe is achieved by Ferreira et al. fer:2024 .
@@
Up to this point, what we need to emphasize is that all the above researches on K-essence are actually one-field K-essence. In specific, the Lagrangian of K-essence is the function of single field $\phi$ and its derivatives $\partial_{\mu}\phi$. However, in this paper, we shall consider two-field K-essence among which there are two scalar fields $\phi$ and $\psi$ together with their derivatives. The novelty of this research is that the Lagrangian functions are significantly different from the previous ones. Exactly, they are constructed from the quotient form“potenials”, $x/y$ with $x$ and $y$ the kinetic term of the fields. To our knowledge, one has never make an investigation on this sort of K-essence. On the other hand, what we consider is a pure K-essence which makes the calculation considerable simple. In fact, we can get the energy density, the pressure and the equation of state (EOS) straightforward because the equations of motion of K-essence can be exactly solved without the help of dynamical analysis. Although it seems our two-field pure K-essence can be transformed to the multi-scalar tensor theories, the equivalence is merely mathematically, not physically at least from the standpoint of speed of perturbations. The reason is that the resulting metric of field-space is not the function of dynamical field, but the auxiliary scalar fields. It is not the case for conventional multi-scalar tensor theories langlois:2008 .
@@
The paper is organized as follows. In Section II, we derive the Einstein equations and the equations of motion for the two-field pure K-essence. In Section III, we study the cosmic evolution of the K-essence in the background of Friedmann-Robertson-Walker Universe. In the first place, the Lagrangian functions of K-essence for some conventional EOS are reconstructed. Then motivated by the Lagrangian of K-essence form for the conventional EOS, we explore the so-called novel K-essence Lagrangian which has the quotient “potentials”. In this way, we obtain several unconventional EQS of K-essence. In Section IV, we transform the two-field pure K-essence to the multi-field form with resort to auxiliary scalar fields. We point that the two frameworks are merely equivalent to each other mathematically, not physically, at least from the point of view of speed of perturbations. In Section V, as the demonstration again for the simplicity of novel two-field pure K-essence, we give a black hole solution. Finally, Section VI gives the conclusion and discussion. Throughout this paper, we adopt the system of units in which $G=c=\hbar=1$ and the metric signature (,+,+,+)(-,+,+,+)( - , + , + , + ).
@@
$x\equiv-\frac{1}{2}\nabla_{\mu}\phi\nabla^{\mu}\phi\;,\ \ \ y\equiv-\frac{1}{2}\nabla_{\mu}\psi\nabla^{\mu}\psi\;.$ (3)
$R$ is the Ricci scalar and $L_{m}$ is the Lagrangian for matters. Make variation of the action with respect to the metric tensor, we obtain the Einstein equations
@@
where $a(t)$ is the scale factor of the Universe. The Einstein equations appears as the Friedmann equation and the acceleration equation
@@ III.2 Three conventional equation-of-states
The so-called three conventional EOSs (equation-of-states) are for the Einstein cosmological constant, the constant EOS and the generalized Chaplygin gas.
@@
with $s$ a constant. This has been pointed by us in Ref. gao:2022 . It is also straightforward to obtain the Schwarzschild-de Sitter solution
@@ III.2.3 Generalized Chaplygin gas
where $A$ and $n$ are positive constants. Observations constrains $n>1$ in order to ensure the condition of stability and the condition of causality.
@@
III.3 several unconventional EOSs
@@ III.3.1 The case of $F=2y+s{\frac{\sqrt{x}}{y^{2n}}}$
In Fig. (1) we plot the equation of state $w$ with respect to the scale factor. In order that the energy density is always positive, we assume the two integration constants $b_{0}>0$ and $b_{1}>0$. The figure shows that when $n=-1/4$, the K-essence behaves as the stiff matter which has the equation of state $+1$. However, it is amazing that its Lagrangian function $F=2y+s\sqrt{xy}$ is obviously different from that for conventional stiff matter $F=-2y$. If $n=1/4$, the K-essence can play the role of cosmic dark energy as mentioned earlier. On the other hand, if $|n|\rightarrow\infty$, the K-essence unifies the Einstein cosmological constant and the cold dark matter model. Finally, when $n\approx{0}$, the equation of state can be very large such that $w\rightarrow{+\infty}$. Now the K-essence becomes the stiffest matter in nature.
@@ III.3.2 The case of $F=s_{1}\sqrt{x}+s_{2}\sqrt{y}+s\left(\frac{{x}}{y}\right)^{\frac{n}{2}}$
In Fig. (2) we plot the EOS $w$ with respect to the scale factor. It shows that the EOS starts from $-1$ at very early universe and then decreases to negative. Subsequently it approaches $-1$ in the distant future. This shows the energy density emerges as nearly a constant at very early universe. Then it gets a remarkable growth. Eventually, the energy density climbs to a very huge constant again by the order of $120$ which is the order Planck energy density. To show this point, we plot the evolution for the Log-ratio
@@
Fig. (6) tells us the early density is the order of $120$ compared to the present-day cosmological constant. So the cosmological constant problem disappears. Fig. (7) shows the EOS can be very large which indicates the K-essence energy density undergoes a precipitous decline at this epoch. Fig. (8) clearly shows the EOS has $-1$ as its starting point and finally consequently result.
@@ III.3.4 The case of $F=s_{1}\sqrt{x}+s_{2}\sqrt{y}+s\frac{\sqrt{x}}{y^{\frac{n}{2}}}$
which is also shown in Fig. (12). Fig. (13), Fig. (14) and Fig. (15) reveal that the EOS can cross the phantom divide either from large to small as shown in Fig. (13) and Fig. (14) or from small to large as shown in Fig. (14). They have one thing in common. Namely, they all approaches $-1$ in distant future.
@@
It is of particular interest when $n=1$. We find that in this case, the K-essence can unify inflation field, dark matter and dark energy. Eq. (47) and Eq. (48) tell us the EOS is $-1$ both at very early universe and later universe which is the desired feature to be the candidate of inflation field and dark energy. In Fig. (16) we plot the EOS of K-essence with respect to the natural logarithm of scale factor when $n=1$. The figure shows inflation starts from $a\sim e^{-140}$ to $a\sim e^{-80}$, i.e., $60$ e-folds are generated by inflation. Then from $a\sim e^{-80}$ to $a\sim e^{-0}$, K-essence plays the role of dark matter with EOS of 00. Finally, at the vicinity of $a^{0}$, namely, the present-day Universe, K-essence plays the role of both dark matter and dark energy. To illustrate this point, we plot the evolution of EOS with respect to cosmic redshift in Fig. (17). The circled line is for $\Lambda CDM$ paradigm. As we can see, the two curves are almost overlap with each other, suggesting that the unification model holds here. In Fig. (18), we plot the evolution of for the Log-ratio $\eta$ of K-essence density to its present-day value with respect to the natural logarithm of scale factor. There is a slope between two platforms. The slope denotes the decreasing of dark matter with the expansion of the Universe.
@@
We recognize immediately that it belongs to the generalized multi-field theories langlois:2008 where the linear perturbations are already completed. Then we conclude from Ref. langlois:2008 that both the entropy modes and adiabatic mode propagate with the speed of light in the framework of Lagrangian function Eq. (50). Now there is an important issue that we need to direct our attention to. The issue is that the speed of perturbations depend on the framework or gauge for the theory. Namely, the speed of perturbations in the framework of Eq. (49) is different from that in the framework of Eq. (50). We can understand this point from a simple example. For simplicity, we consider one-field pure K-essence with the Lagrangian
@@
In conclusion, we have studied a novel two-field pure K-essence whose Lagrangian is the function of kinetic terms of two scalar fields while independent on the scalar fields. The absence of the scalar fields variables in the Lagrangian makes the equations of motion for scalar fields considerable simple and their analytic solutions can be found. The novelty of this k-essence lies in the ‘potential” which is made up with quotient of kinetic terms of the scalar fields.
@@
4. Since the EOS begins from $-1$ and ends in $-1$, K-essence unifies the early inflation field, the later dark energy and the dark matter in between.
@@
6. The equivalence of this novel K-essence to the multi-field theory is merely from the perspective of mathematics, not physics. The reason is that the metric of field space is not the function of dynamical scalar fields, but the function of auxiliary scalar fields.

### 12 Feb 2024 v2

@@ I Introduction
K-essence is a certain class of scalar field models with nonstandard kinetic terms. The name was first coined in a paradigm for inflation by Armendariz-Picon et al picon:1999 ; garriga:1999 . Subsequently, it was found that k-essence can present interesting models for dark energy chiba:2000 ; picon:2000 ; picon:2001 ; chiba:2002 ; chimento:2004 ; chimento:2003 ; malquarti:2003 . Right after these studies, Scherrer shows that K-essence can also serve as the unified model for dark matter and dark energy robert:2004 . Afterwards, researches on the K-essence theories spring up.
@@
Specifically, Chimonto and Lazkoz make an analysis on the atypical K-essence cosmologies chimento:2005 . Abramo et al conduct the research on the stability problem of phantom k-essence theories abramo:2006 . Bonvin et al bonvin:2006 demonstrate that if one uses K-essence to solve the coincidence problem and play the role of dark energy in the Universe, the fluctuations of the field have to propagate superluminally at some stage. This implies that successful K-essence models violate causality. Therefore, K-essence cannot arise as a low energy effective field theory of a causal, consistent high energy theory. Bamba bamba:2012 et al carry out a thorough investigation on the cosmological perturbations of K-essence. The causality and superluminal behavior K-essence theories are investigated by Bruneton jean:2007 . The slow-roll conditions for thawing K-essence are derived by Chiba et al with a separable Lagrangian K-essence chiba:2009 . The K-essence inflation model in the framework of $f(R)$ gravity is studied by Nojiri et al nojiri:2019 . The phenomenological implications of the constant-roll condition for the K-essence inflation is investigated by Odintsov and Oikonomou odintsov:2020 . The late-time behavior of a pure K-Essence in $f(R)$ gravity is studied by Odintsov et al odintsov:2020pdu . Considering the unification between the swampland criteria and the attractor, Herrera studies the reconstruction of the K-essence inflationary universe herrera:2020 . Gialamasa and Lahanasa show that the reheating in $R^{2}$ Palatini gravity in the Einstein frame, shares common features with K-essence models of inflation gia:2020 . The inflationary phenomenology of a K-essence inflation in the framwork of Einstein-Gauss-Bonnet gravity is conducted by Odintsov et al odintsov:2021 . Furthermore, K-essence can yield a complete picture of the evolution of the Universe, starting from the early inflation, the subsequent dark matter domination, and to the late time cosmic acceleration bose:2009a ; bose:2009b . Interestingly, Tian and Zhu’s research tian:2021 reveals that K-essence framework is also conducive to the Hubble tension problem riess:2021 ; freedman:2021 . [^2] But Lee et al lee:2022 prove that the EOS (equation of state) of dark energy and the present-day Hubble parameter are anti-correlated in K-essence theories. So K-essence as latetime dark energy may make the Hubble tension even worse. The reason is that if the K-eesence scalar does not rest at the minimum of the potential, the present-day Hubble parameter would be decreased.
@@
Very recently, Hung and Miao hung:2023 investigate whether a quantum-induced K-essence can provide efficient reheating without affecting the observational constraints on primordial inflation. Kehayias and Scherrer find a new generic evolution for K-essence dark energy with the equation of state of roughly cosmological constant kehayias:2019 . Sarkar et al sarkar:2023 show that the accretion of k-essence dark energy can lead to a faster circularising of orbits for binary black holes. Lara et al lara:2022 investigate in depth the relation between K-essence, well-posedness of the Cauchy problem and the UV completions. Huang huang:2021 constructs the thawing K-essence models by generating Taylor expansion coefficients of Lagrangian functions from random matrices. Chatterjee et al make a study on the dynamical stability problem for K-essence interacting nonminimally with a perfect fluid chatt:2021 . Applying the “$2+1+1$” decomposition of spacetime, Nagy et al conduct the research of spherically symmetric, static black holes with scalar hair, and naked singularities in nonminimally coupled K-essence nagy:2021 . Barvinsky et al find that the generalized unimodular gravity can be transformed into a new form of K-essence barvinsky:2021 . Finally, the K-essence model in which K-essence can be responsible for both primordial inflation and the present observed acceleration, while also admitting a nonsingular de Sitter beginning of the Universe is achieved by Ferreira et al fer:2024 .
@@
Up to this point, what we need to emphasize is that all the above researches on K-essence are actually one-field K-essence. In specific, the Lagrangian of K-essence is the function of single field $\phi$ and its derivatives $\partial_{\mu}\phi$. However, in this paper, we shall consider two-field K-essence among which there are two scalar fields $\phi$ and $\psi$ together with their derivatives. The novelty of this research is that the Lagrangian functions are significantly different from the previous ones. Exactly, they are constructed from the quotient form“potenials”, $x/y$ with $x$ and $y$ the kinetic term of the fields. To our knowledge, one has never make an investigation on this sort of K-essence. On the other hand, what we consider is a pure K-essence which makes the calculations considerable simple. In fact, we can get the energy density, the pressure and the equation of state straightforward because the equations of motion of K-essence can be exactly solved without the help of dynamical analysis. Although it seems our two-field pure K-essence can be transformed to the multi-scalar tensor theories, the equivalence is merely mathematically, not physically at least from the standpoint of speed of perturbations. The reason is that the resulting metric of field-space is not the function of dynamical field, but the auxiliary scalar fields. It is not the case for conventional multi-scalar tensor theories langlois:2008 .
@@
The paper is organized as follows. In Section II, we derive the Einstein equations and the equations of motion for the two-field pure K-essence. In Section III, we study the cosmic evolution of the K-essence in the background of Friedmann-Robertson-Walker Universe. In the first place, the Lagrangian functions of K-essence for some conventional EOS are reconstructed. Then motivated by the Lagrangian of K-essence form for the conventional EOS, we explore the so-called novel K-essence Lagrangian which has the quotient “potentials”. In this way, we obtain several unconventional EOSs (equation-of-states) of K-essence. In Section IV, we transform the two-field pure K-essence to the multi-field form with resort to auxiliary scalar fields. We point out that the two frameworks are merely equivalent to each other mathematically, not physically, at least from the point of view of speed of perturbations. In Section V, as the demonstration again for the simplicity of novel two-field pure K-essence, we give a black hole solution. Finally, Section VI gives the conclusion and discussion. Throughout this paper, we adopt the system of units in which $G=c=\hbar=1$ and the metric signature (,+,+,+)(-,+,+,+)( - , + , + , + ).
@@
$x\equiv-\frac{1}{2}\nabla_{\mu}\phi\nabla^{\mu}\phi\;,\ \ \ y\equiv-\frac{1}{2}\nabla_{\mu}\psi\nabla^{\mu}\psi\;,$ (3)
and they are positive with the signature of (,+,+,+)(-,+,+,+)( - , + , + , + ) in the Friedamnn-Roberstson-Walker Universe. $R$ is the Ricci scalar and $L_{m}$ is the Lagrangian for matters. Make variation of the action with respect to the metric tensor, we obtain the Einstein equations
@@
where $a(t)$ is the scale factor of the Universe. The Einstein equations appear as the Friedmann equation and the acceleration equation
@@ III.2 Three conventional equation-of-states
The so-called three conventional EOSs are for the Einstein cosmological constant, the constant EOS and the generalized Chaplygin gas.
@@
with $s$ a constant. This has been pointed out by us in Ref. gao:2022 . It is then straightforward to obtain the Schwarzschild-de Sitter solution
@@ III.2.3 Generalized Chaplygin gas
where $A$ and $n$ are positive constants. Observations constrain $n>1$ in order to ensure the condition of stability and the condition of causality.
@@
III.3 four unconventional EOSs
@@ III.3.1 The case of $F=2y+s{\frac{\sqrt{x}}{y^{2n}}}$
In Fig. (1) we plot the equation of state $\omega$ with respect to the scale factor. In order that the energy density is always positive, we assume the two integration constants $b_{0}>0$ and $b_{1}>0$. The figure shows that when $n=-1/4$, the K-essence behaves as the stiff matter which has the equation of state $+1$. However, it is amazing that its Lagrangian function $F=2y+s\sqrt{xy}$ is obviously different from that for conventional stiff matter $F=-2y$. If $n=1/4$, the K-essence can play the role of cosmic dark energy as mentioned above. On the other hand, if $|n|\rightarrow\infty$, the K-essence unifies the Einstein cosmological constant and the cold dark matter model. Finally, when $n\approx{0}$, the equation of state can be very large such that $w\rightarrow{+\infty}$. Now the K-essence becomes the stiffest matter in nature.
@@ III.3.2 The case of $F=s_{1}\sqrt{x}+s_{2}\sqrt{y}+s\left(\frac{{x}}{y}\right)^{\frac{n}{2}}$
In Fig. (2) we plot the EOS $\omega$ with respect to the scale factor. It shows that the EOS starts from $-1$ at very early universe and then decreases to negative. Subsequently it approaches $-1$ in the distant future. This shows the energy density emerges as nearly a constant at very early universe. Then it gets a remarkable growth. Eventually, the energy density climbs to a very huge constant again by the order of $120$ which is the order Planck energy density. To show this point, we plot the evolution for the Log-ratio
@@
Fig. (6) tells us the early density is the order of $120$ compared to the present-day cosmological constant. So the cosmological constant problem disappears. Fig. (7) shows the EOS can be very large which indicates the K-essence energy density undergoes a precipitous decline at this instant. Fig. (8) clearly shows the EOS has $-1$ as its starting point and finally consequently result.
@@ III.3.4 The case of $F=s_{1}\sqrt{x}+s_{2}\sqrt{y}+s\frac{\sqrt{x}}{y^{\frac{n}{2}}}$
which is also shown in Fig. (12). Fig. (13), Fig. (14) and Fig. (15) reveal that the EOS can cross the phantom divide either from large to small as shown in Fig. (13) or from small to large as shown in Fig. (14). They have one thing in common. Namely, they all approaches $-1$ in distant future.
@@
It is of particular interest when $n=1$. We find that in this case, the K-essence can unify inflation field, dark matter and dark energy. Eq. (47) and Eq. (48) tell us the EOS is $-1$ both at very early universe and later universe which is the desired feature to be the candidate of inflation field and dark energy. In Fig. (16) we plot the EOS of K-essence with respect to the natural logarithm of scale factor when $n=1$. The figure shows inflation starts from $a\sim e^{-140}$ to $a\sim e^{-80}$, i.e., $60$ e-folds are generated by inflation. Then from $a\sim e^{-80}$ to $a\sim e^{-0}$, K-essence plays the role of dark matter with EOS of 00. Finally, at the vicinity of $a^{0}$, namely, the present-day Universe, K-essence plays the role of both dark matter and dark energy. To illustrate this point, we plot the evolution of EOS with respect to cosmic redshift in Fig. (17). The circled line is for $\Lambda CDM$ paradigm. As we can see, the two curves are almost overlap with each other, suggesting that the unification model holds here. In Fig. (18), we plot the evolution for the Log-ratio $\eta$ of K-essence density to its present-day value with respect to the natural logarithm of scale factor. There is a slope between two platforms. The slope denotes the decrease of dark matter with the expansion of the Universe.
@@
We recognize immediately that it seems belonging to the generalized multi-field theories langlois:2008 where the linear perturbations are already completed. Then we conclude from Ref. langlois:2008 that both the entropy modes and adiabatic mode propagate with the speed of light in the framework of Lagrangian function Eq. (50). Now there is an important issue that we need to direct our attention to. The issue is that the speed of perturbations depend on the framework or gauge for the theory. Namely, the speed of perturbations in the framework of Eq. (49) is different from that in the framework of Eq. (50). We can understand this point from a simple example. For simplicity, we consider one-field pure K-essence with the Lagrangian
@@
In conclusion, we have studied a novel two-field pure K-essence whose Lagrangian is the function of kinetic terms of two scalar fields while independent on the scalar fields. The absence of the scalar fields variables in the Lagrangian makes the equations of motion for scalar fields considerable simple and their analytic solutions can be found. The novelty of this K-essence lies in the “potential” which is made up with quotient of kinetic terms of the scalar fields.
@@
4. Since the EOS begins from $-1$, goes through 00 and ends in $-1$, K-essence unifies the early inflation field, the later dark energy and the dark matter in between.
@@
6. The equivalence of this novel K-essence to the multi-field theory is merely from the perspective of mathematics, not physics. The reason is that the metric of field-space is not the function of dynamical scalar fields, but the function of auxiliary scalar fields.